Solve the following system of equations for z and for y System of Equations: Value of z Value of y y 65- Sz Solve the following system of equations for a and for b: System of Equations: 4а + 2b-: 20 5a +5b 60 Value of a Value of b Plot the following system of equations on the following graph System of Equations: p- 6-2q Note: Use the orange line (square symbols) to plot the first equation, and use the blue line (circle symbols) to plot the second equation. 10 0 207 2
Note: Use the orange line (square symbols) to plot the first equation, and use the blue line (aircle symbols) to plot the second equation. 10 0 0 123567 80 The values of p and q that solve these two equations simultaneously can be seen on the graph as Grade It Now Save& Continue Continue without saving
Answer #1

y = 9 + 3z

y - 3z = 9

y = 65 - 5z

y + 5z = 65 ( 1)

y - 3z = 9 ( 2)

Subtracting equating 1 and 2

8z = 56

Z = 7

y = 9 + 3z

y = 9 + 3 x 7

y = 30

Value of z = 7

Value of Y = 30

_____________________________________________________________________

4a + 2b = 20 ( 1)

15a + 5b = 60 ( 2)

Multiplying equating 1 by 3.75

15a + 7.5b = 75 (3)

15a + 5b = 60 (4)

Subtracting equating 3 from 4

we get

2.5b = 15

b = 6

4a + 2b = 20

4a = 20 - 2b

4a = 20 - 2 x 6

4a = 8

a = 2

Value of a = 2

Value of b = 6

_____________________________________________________________________________

p = 6 - 2q

3 2 1 0 76543.210

p = 4 + q

25 20 15 10 5 0 5 15 20